OpenStudy (anonymous):

What is the 32nd term of the arithmetic sequence where a1 = 12 and a13 = -60?

OpenStudy (campbell_st):

a term in an arithmetic series can be found using \[T_{n} = a_{1} + (n - 1)d\] this can be used with the a13 and a1 to find the common difference (d) \[-60 = 12 + (13 - 1)d ,,,, or ... -60 = 12 + 12d\] solve for d when you have d, you can find the 32 term by \[T_{32} = 12 + (32 -1)d\]

OpenStudy (anonymous):

So the answer would be -174?

OpenStudy (campbell_st):

what was d..?

OpenStudy (anonymous):

-6

OpenStudy (campbell_st):

yep... so the answer is correct... well done.

OpenStudy (anonymous):

Thanks!